# Bias-Field Digitized Counterdiabatic Quantum Algorithm for Higher-Order Binary Optimization

### v1

@@ III.1 HUBO instances considered
Figure 1: 433-qubit NN HUBO results on IBM Osprey simulated via MPS. (a) Energy distributions for iterations 0, 2, 4, 6, 8 and 10; from grey to maroon using our novel BF-DCQO variant. The continuous line corresponds to the reference ground-state energy $E_{0}$ given by DMRG, and the dashed lines correspond to the minima obtained from the distributions. (b) Same plot as (a), going from grey to teal, but considering the original BF-DCQO approach [48]. (c)-(d) Approximation ratios and distances to solution obtained for each iteration for both approaches. For both panels and only for our variant (maroon markers), the shaded area indicates that the bias update is changed from unsigned to signed bias in the last iteration. See Table 1.
Table 1: Resources needed and results obtained per BF-DCQO iteration to optimize via MPS a 433-qubit NN HUBO with $n_{\text{shots}}=10000$ and $\theta_{\text{cutoff}}=0.06$ for both our BF-DCQO variant (left columns) and the original approach (right columns). For our variant, we used $\alpha=0.01$ and we perform a signed bias in the last iteration.
@@
0 497 497 5910 5910 5934 5934 2275 2275 99 99 0.480 0.480 0.358 0.366
1 470 453 5563 5421 5525 5379 2156 2108 98 98 0.615 0.592 0.225 0.285
2 466 446 5412 5310 5365 5247 2106 2066 98 98 0.689 0.632 0.158 0.247
3 436 427 5314 5314 5254 5261 2063 2065 98 98 0.724 0.654 0.148 0.214
4 455 441 5283 5310 5240 5242 2052 2063 98 98 0.745 0.681 0.118 0.173
5 479 450 5287 5310 5238 5251 2052 2063 98 98 0.760 0.705 0.102 0.175
6 474 448 5287 5301 5247 5207 2052 2061 99 98 0.769 0.709 0.096 0.167
7 455 435 5283 5301 5198 5226 2052 2061 98 99 0.776 0.712 0.089 0.156
8 464 456 5283 5310 5210 5251 2052 2063 98 98 0.781 0.705 0.074 0.156
9 447 447 5284 5277 5210 5214 2052 2052 99 98 0.785 0.709 0.064 0.162
10 100 456 1879 5283 1653 5225 496 2052 80 97 0.954 0.710 0.019 0.146
In contrast to physical phase transitions where variations of the system parameters lead to qualitative changes in its properties, in computer science, computational phase transitions occur at critical points where algorithms require an increasing amount of computational resources, becoming less and less tractable [76]. The computational phase transitions of $k$-SAT have been the subject of several works [76, 77, 78]. They depend critically on the density of the problem, i.e., the ratio between its clauses and variables.
III.2 Comparison of methods
@@
In the following lines, we perform a comparison of both the original BF-DCQO approach [48] and our novel proposed variant, conducting a 433-qubit study using a NN HUBO instance [Eq. (12)] via MPS with which to simulate the expected performance of the upcoming IBM Osprey quantum platform without noise. For this case, the values $n_{\text{shots}}=10000$ and $\theta_{\text{cutoff}}=0.06$ are set for both approaches and $\alpha=0.01$ for our variant. The results can be seen in Fig. 1 with the amount of resources needed per iteration in Table 1.
@@
Figure 1 shows that before applying the signed bias, we obtain a $10.6\%$ enhancement of the AR and $60.8\%$ for the DS, demonstrating that our novel CVaR-based approach for updating the bias fields provides a better performance than standard BF-DCQO. After the signed bias method is applied for the last iteration, these values increases to $34.4\%$ and $87.3\%$, respectively, showing that this additional feature also remarkably contributes to improving the results obtained from standard BF-DCQO.
@@
In addition, we find as an intrinsic signature of our BF-DCQO variant that the circuits implemented for each iteration require less and less resources even though the outcomes are more optimal. Conceptually, the goal of BF-DCQO is to encode the ground state of $H_{f}$, which is a product state in the computational basis. As this is a separable state with no entanglement, BF-DCQO pushes the energy distribution towards the preparation of this state, thereby reducing the amount of entanglement required per iteration.
@@
Figure 2: 156-qubit NN HUBO instance described by Eq. (12) under different approaches: simulated annealing (SA, blue), Tabu search (TS, orange), D-Wave (green), BF-DCQO om ibm_fez (purple), and the best outcome of QAOA up to $p=3$ layers (dashed red). The DMRG optimal solution is used as a ground state reference (black line). The dashed vertical lines indicate the minimum energy values obtained for each approach. The D-Wave and BF-DCQO distributions correspond to experiments run on quantum hardware.
@@
  • Simulated annealing (SA) and Tabu search (TS) both with $n_{\text{shots}}=100000$.
@@
  • Tenth iteration of BF-DCQO [48], with $\alpha=0.02$ run on ibm_fez using $\theta_{\text{cutoff}}=0.06$, $n_{\text{trot}}=3$ and $n_{\text{shots}}=10000$. Previously, the ten BF-DCQO iterations were simulated using MPS without noise.
@@
  • The DMRG implementation of the ITensor library [74] setting 5 sweeps, maximum bond dimensions of $[5,10,10,10,20]$ and a truncation error cutoff of $10^{-5}$ (see Appendix B). We sample over 100 random MPS initial states with bond dimension $\chi=5$ and select the best outcome.
Figure 2 illustrates that BF-DCQO is capable of outperforming other established methods on current noisy quantum platforms, returning an energy distribution whose minimum outcome is very close to the ground state reference provided by DMRG. In particular, we have obtained a $26.7\%$ enhancement of the AR and $92.7\%$ of the DS metrics with respect to D-Wave and a $72.8\%$ enhancement of DS for SA approach but a small $-8.0\%$ lowering for the AR metric. Despite the fact that three layers were chosen for CVaR-QAOA to recover the expected behaviour of BF-DCQO with $n_{\text{trot}}=3$, the obtained results show the worst performance, even in comparison with D-Wave. This finding is in accordance with the results presented in Ref. [40]. Furthermore, it should be noted that BF-DCQO and QAOA are the only approaches that do not suffer from the inherent qubit overhead required by the rest of the methods considered for mapping HUBO into QUBO problems. In addition, as the order of many-body terms in the problem Hamiltonian $H_{f}$ increases, the qubit overhead increases and more resources are required. In particular, regular quadratization methods require a number of ancillary qubits proportional to $n_{k\text{-body}}(k-2)$, with $n_{k\text{-body}}$ the number of $k$-body terms of $H_{f}$ [79, 80, 81]. Therefore, a worse performance is expected. Other known methods such as Gurobi [82] and CPLEX [83] also suffer from this drawback. For the results presented in Fig. 2, 334 variables were needed for SA, TS and D-Wave. This feature is absent in BF-DCQO, making it a suitable technique for working with large-scale optimization problems with minimal resource requirements.
Up to this point, we have compared the performance of BF-DCQO on quantum hardware using only the last iteration of the entire routine simulated via MPS. In the following section, the impact of running the entire routine under noisy environments is examined. To this end, larger values of $\theta_{\text{cutoff}}$ will be set in order to guarantee implementability. We note that considering a larger number of measurements might allow us to partially recover the lost expressivity due to the increase of $\theta_{\text{cutoff}}$.
III.3 Hardware implementation on IBM
We first make use of the 156-qubit superconducting quantum platform ibm_fez with the native gate set $\{X,\sqrt{X},\text{RZ}(\theta),\text{CZ}\}$ (see Appendix C). The linearity of its heavy-hex qubit coupling map encourages us to work with Hamiltonians whose many-body terms are nearest neighbours, thereby avoiding an overhead of entangling gates for coupling distant qubits. With this in mind, we solve the same 156-qubit NN HUBO instance that was previously tested in Fig. 2, but we now run the entire routine on ibm_fez. The computation is thus conducted in a purely quantum manner. We use $n_{\text{shots}}=10000$, $\alpha=0.01$, $\theta_{\text{cutoff}}=0.13$ and a total of eleven iterations, considering an unsigned-bias update for the first ten and a signed bias for the last one. The results are presented in Fig. 3, which depicts the iterative procedure of BF-DCQO and how the obtained outcomes move towards the desired solution. The amount of resources needed per iteration is given in Table 2. As a final remark, the final iteration of the purely quantum BF-DCQO routine exhibits a slightly worse performance when compared to the results of Fig. 2, which employed a noiseless simulation with a considerably lower gate threshold. The purely quantum BF-DCQO still outperforms the optimal solution obtained on the D-Wave platform, where we obtained $34.1\%$ of AR enhancement and $66.4\%$ of DS enhancement with respect the experiment run on D-Wave in Fig. 2.
@@ III.3 Hardware implementation on IBM
Figure 3: 156-qubit NN HUBO results on ibm_fez. (a) Energy distributions for iterations 0, 2, 4, 6, 8 and 10; from grey to maroon respectively. The continuous line corresponds to the reference ground state $E_{0}$ given by DMRG. The dashed lines correspond to the minima obtained from the distributions. (b)-(c) Approximation ratios and distances to solution for each iteration. For both panels, the shaded area indicates that the bias update is changed from unsigned to signed bias in the last iteration. See Table 2.
@@
Table 2: Resources needed and results obtained per BF-DCQO iteration on ibm_fez to optimize a 156-qubit HUBO with $n_{\text{shots}}=10000$, $\alpha=0.01$ and $\theta_{\text{cutoff}}=0.13$. At the last iteration, we performed signed bias.
@@
10 3 330 180 13 39 0.815 0.037
Figure 4: 156-qubit MAX W-3-SAT HUBO results on ibm_fez. (a) Energy distributions for iterations 0, 2, 4, 6, 8 and 10; from grey to maroon respectively. The continuous line corresponds to the reference ground-state energy $E_{0}$ given by DMRG. The dashed lines correspond to the maxima of the distributions. (b)-(c) Approximation ratios and distances to solution obtained for each iteration. The shaded area indicates that the bias update is changed from unsigned to signed in the last iteration. See Table 3.
Table 3: Resources needed and results obtained per BF-DCQO iteration on ibm_fez to optimize a 156-qubit NN MAX W-3-SAT HUBO problem with $n_{\text{shots}}=10000$, $\alpha=0.01$, and $\theta_{\text{cutoff}}=0.06$. In the last iteration, we apply a signed bias, where no entangling gates are used.
@@
With the aim of solving more industrially relevant and denser large-scale problems, we also solve a 156-qubit NN MAX W-3-SAT instance on ibm_fez, described by Eq. (15) with only nearest-neighbour terms. We use the same parameters as for the previous instance but change $\theta_{\text{cutoff}}=0.06$ for convenience. The results are shown in Fig. 4, and the amount of resources needed per iteration is listed in Table 3. For this particular problem, the best outcome obtained is remarkably close to the optimal solution despite a purely quantum routine, which highlights the efficacy of BF-DCQO under noisy environments. Additionally, we solved the same HUBO instance on D-Wave using $n_{\text{shots}}=100000$, where again the purely quantum BF-DCQO outperforms their results, where we obtained $43.0\%$ of AR enhancement and $88.5\%$ of DS enhancement with respect the experiment run on the D-Wave platform.
IV Conclusion

### v2

@@ Methods
In contrast to physical phase transitions where variations of the system parameters lead to qualitative changes in its properties, in computer science, computational phase transitions occur at critical points where algorithms require an increasing amount of computational resources, becoming less and less tractable [77]. The computational phase transitions of $k$-SAT have been the subject of several works [77, 78, 79]. They depend critically on the density of the problem, i.e., the ratio between its clauses and variables. For random 3-SAT instances, it has been proven that the critical point is located around a density of $\alpha_{C}=4.3$.
Figure 2: Performance comparison of the BF-DCQO variants for the 433-qubit NN HUBO results simulated via MPS. a, Energy distributions for iterations 1, 3, 5, 7, 9 and 11; from grey to maroon using our novel BF-DCQO variant. The continuous line corresponds to the exact ground-state energy $E_{0}$ given by an iterative method [Supplementary Note 2], and the dashed lines correspond to the minima obtained from the distributions. Here, we report energies in units of the standard deviation of the coupling constants in Eq. (13). b, Same plot as a, going from grey to teal, but considering the original BF-DCQO approach [51]. c-d, Approximation ratios and distances to solution obtained for each iteration for both approaches. For both panels and only for our variant (maroon markers), the shaded area indicates that the bias update is changed from unsigned to signed bias in the last iteration. See Table 1.
Table 1: 433-qubit NN HUBO resources needed and results simulated via MPS. We compare both our BF-DCQO variant (left columns) and the original approach (right columns) with $n_{\text{shots}}=10000$ and $\theta_{\text{cutoff}}=0.06$. For our variant, we used $\alpha=0.01$ and we perform a signed bias in the last iteration.
@@
1 497 497 5910 5910 5934 5934 2275 2275 99 99 0.477 0.477 0.361 0.370
2 470 453 5563 5421 5525 5379 2156 2108 98 98 0.612 0.589 0.229 0.289
3 466 446 5412 5310 5365 5247 2106 2066 98 98 0.686 0.628 0.163 0.251
4 436 427 5314 5314 5254 5261 2063 2065 98 98 0.720 0.651 0.152 0.218
5 455 441 5283 5310 5240 5242 2052 2063 98 98 0.741 0.678 0.123 0.178
6 479 450 5287 5310 5238 5251 2052 2063 98 98 0.756 0.702 0.106 0.180
7 474 448 5287 5301 5247 5207 2052 2061 99 98 0.765 0.705 0.101 0.171
8 455 435 5283 5301 5198 5226 2052 2061 98 99 0.772 0.708 0.094 0.161
9 464 456 5283 5310 5210 5251 2052 2063 98 98 0.777 0.701 0.078 0.161
10 447 447 5284 5277 5210 5214 2052 2052 99 98 0.781 0.706 0.068 0.166
11 100 456 1879 5283 1653 5225 496 2052 80 97 0.949 0.706 0.023 0.150
von Neumann entanglement entropy
The role of entanglement in quantum optimization is an interesting question that has been discussed recently in the context of adiabatic [80, 81] and variational [82, 83, 84, 85, 86] algorithms, with no clear consensus. Here, we study how entanglement changes with each bias-field iteration, by means of MPS simulations, to support our understanding of how BF-DCQO converges to the final solution.
The entanglement between two subsystems, in the case of a composite pure state, can be quantified by the von Neumann entropy of the reduced density matrix of either subsystem. In an MPS representation of the wavefunction [87], the entanglement entropy for any bipartition is given by the singular values corresponding to the bond that connects the two parts. For bond index $i$, we denote the $j$th-largest singular value by $\lambda_{ij}$. Here, $j$ has values $j\in\{1,2,\dots,\chi_{i}\}$ up to the bond dimension $\chi_{i}$ and the bonds between qubits (tensors in the MPS) are indexed by $i\in\{1,\dots,N-1\}$. We consider the entanglement entropy averaged over bonds
$S=-\frac{1}{N-1}\sum_{i=1}^{N-1}\sum_{j=1}^{\chi_{i}}\lambda^{2}_{ij}\log_{2}\lambda^{2}_{ij}.$ (18)
This quantity is minimally zero when the state is a product state and has an upper bound $\sim\log_{2}2^{N}=N$ when the reduced density matrix of either subsystem is maximally mixed [88]. While entanglement generally scales with the subsystem volume, in special cases, stricter bounds such as area laws can be derived [89, 90]. Nonequilibrium states resulting from global parameter quenches are expected to follow a volume law, though recent quantum simulations of paramagnetic-to-spin-glass quench dynamics exhibited area-law scaling [34]. Here, we do not analyze the scaling behavior but rather focus on the entanglement generated in the short-time dynamics of the counterdiabatically driven systems, as discussed in Entanglement in the iterative optimization.
Results and discussion
Enhanced BF-DCQO
@@
In the following lines, we perform a comparison of the original BF-DCQO approach [51] and our proposed variant. We study a 433-qubit NN HUBO instance [Eq. (13)], using MPS to simulate the expected performance of the upcoming IBM Osprey quantum platform without noise. In this case, the values $n_{\text{shots}}=10000$ and $\theta_{\text{cutoff}}=0.06$ are set for both approaches and $\alpha=0.01$ for our variant. The results can be seen in Fig. 2 and the amount of resources needed per iteration in Table 1.
@@
Figure 2 shows that before applying the signed bias, we obtain a $10.6\%$ enhancement of the AR and $59.1\%$ for the DS, demonstrating that our new CVaR-based approach for updating the bias fields provides a better performance than standard BF-DCQO. After the signed-bias method is applied in the last iteration for our variant, these values increase to $34.4\%$ and $84.4\%$, respectively, showing that this additional feature remarkably contributes to improving the results obtained from standard BF-DCQO.
@@
In addition, we find as an intrinsic signature of our BF-DCQO variant that the circuits implemented for each iteration require less and less resources even while the outcomes are more optimal. Conceptually, the goal of BF-DCQO is to encode the ground state of $H_{f}$, which can be either a product state in the computational basis or a superposition of them if the optimal solution is degenerate. For the first case—a separable state with no entanglement—BF-DCQO pushes the energy distribution towards the preparation of this state, reducing the amount of entanglement required per iteration. This feature indicates that BF-DCQO returns a lower-resource-demanding circuit per iteration, yielding a more classically tractable instance. Therefore, the combination of both quantum and classical computation paradigms appears like a promising route to address optimization problems by means of our BF-DCQO protocol. In the case of a superposition ground state, entanglement might still be reduced depending on the instance. We discuss the reduction of entanglement in more detail in Entanglement in the iterative optimization.
Figure 3: 156-qubit NN HUBO instance under different approaches. We solve Eq. (13) considering simulated annealing (SA, blue), local search algorithm (LS, purple), Tabu search (TS, orange), D-Wave (green), BF-DCQO on ibm_fez (maroon), the best outcome of QAOA up to $p=3$ layers (dashed red), DMRG solution (dashed grey) and the exact solution (black line). The last two mostly overlap as DMRG reaches a nearly optimal solution [Supplementary Note 4].
The dashed vertical lines indicate the minimum energy values obtained for each approach. The D-Wave and BF-DCQO distributions correspond to experiments run on quantum hardware. Energies are in units the standard deviation of the coupling constants in Eq. (13).
Comparison of methods
@@
  • Simulated annealing (SA) [91] with $100000$ number of reads, initial (final) temperature $T_{i}=2$ ($T_{f}=0.05$) using a geometric cooling scheduling with $1000$ sweeps.
  • Local search (LS) algorithm, which can be seen as SA with zero-temperature Metropolis-Hastings criterion [92, 93], with $100000$ reads and $10000$ sweeps.
  • Tabu search (TS) [94] with $100000$ reads starting from randomly generated initial states.
@@
  • Tenth iteration of BF-DCQO [51], with $\alpha=0.02$ run on ibm_fez using $\theta_{\text{cutoff}}=0.06$, $n_{\text{trot}}=3$ and $n_{\text{shots}}=10000$, ten times less than SA, TS and D-Wave to compensate the number of iterations. Previously, the ten BF-DCQO iterations were simulated using MPS without noise.
@@
  • The DMRG implementation of the ITensor library [95]. We find convergence with 10 sweeps, maximum bond dimension $10$, truncation error cutoff of $10^{-10}$, and a noise parameter that is gradually reduced from $10^{-3}$ to zero through the sweeps (see Supplementary Note 4).
  • Exact solution using a numerical iterative method [Supplementary Note 2].
@@
Figure 4: 156-qubit NN HUBO results on ibm_fez. a, Energy distributions for iterations 1, 3, 5, 7, 9 and 11; from grey to maroon respectively. The continuous line corresponds to the exact ground state energy $E_{0}$ [Supplementary Note 2]. The dashed lines correspond to the minima obtained from the distributions. Energies are in units the standard deviation of the coupling constants in Eq. (13). b-c, Approximation ratios and distances to solution for each iteration. For both panels, the shaded area indicates that the bias update is changed from unsigned to signed bias in the last iteration. See Table 2.
Figure 5: 156-qubit MAX W-3-SAT HUBO results on ibm_fez. a, Cost function distributions for iterations 1, 3, 5, 7, 9 and 11; from grey to maroon respectively. The continuous line corresponds to the exact ground-state energy $E_{0}$ [Supplementary Note 2]. The cost function is in units of $\max(\omega_{ijk})$ in Eq. (16). The dashed lines correspond to the maxima of the distributions. b-c, Approximation ratios and distances to solution obtained for each iteration. The shaded area indicates that the bias update is changed from unsigned to signed in the last iteration. See Table 3.
Figure 3 illustrates that BF-DCQO is capable of outperforming other established methods on current noisy quantum platforms, returning an energy distribution whose minimum outcome is very close to the exact ground state energy. In particular, we have obtained a $26.7\%$ enhancement of the AR and $92.3\%$ of the DS metrics with respect to D-Wave and a $71.7\%$ enhancement of DS for SA approach but a small $8.0\%$ lowering for the AR metric. Despite the fact that three layers were chosen for CVaR-QAOA to recover the expected behaviour of BF-DCQO with $n_{\text{trot}}=3$, the obtained results show the worst performance, even in comparison with D-Wave. This finding is in accordance with the results presented in Ref. [42]. Furthermore, it should be noted that BF-DCQO and QAOA do not suffer from the inherent qubit overhead required for mapping HUBO into QUBO problems, as D-Wave does. In addition, as the order of many-body terms in the problem Hamiltonian $H_{f}$ increases, the qubit overhead increases and more resources are required. In particular, regular quadratization methods require a number of ancillary qubits proportional to $n_{k\text{-body}}(k-2)$, with $n_{k\text{-body}}$ the number of $k$-body terms of $H_{f}$ [96, 97, 98]. Therefore, a worse performance is expected. Other known methods such as Gurobi [99] and CPLEX [100] also suffer from this drawback. For the results presented in Fig. 3, 334 variables were needed for D-Wave. This feature is absent in BF-DCQO, making it a suitable technique for working with large-scale optimization problems with minimal resource requirements.
Up to this point, we have compared the performance of BF-DCQO to other methods by simulating the entire BF-DCQO routine with MPS and running the circuit obtained in the last iteration on hardware. In the following section, the impact of running the entire routine under noisy environments is examined. To this end, larger values of $\theta_{\text{cutoff}}$ will be set in order to guarantee implementability. We note that considering a larger number of measurements might allow us to partially recover the lost expressivity due to the increase of $\theta_{\text{cutoff}}$.
Hardware implementation on IBM
We first make use of the 156-qubit superconducting quantum platform ibm_fez with the native gate set $\{X,\sqrt{X},\text{RZ}(\theta),\text{CZ}\}$ (see Supplementary Note 3). The linearity of its heavy-hex qubit coupling map encourages us to work with Hamiltonians whose many-body terms are nearest neighbours, thereby avoiding an overhead of entangling gates for coupling distant qubits. With this in mind, we solve the same 156-qubit NN HUBO instance that was previously tested in Fig. 3, but we now run the entire routine on ibm_fez. The computation is thus conducted in a purely quantum manner. We use $n_{\text{shots}}=10000$, $\alpha=0.01$, $\theta_{\text{cutoff}}=0.13$ and a total of eleven iterations, considering an unsigned-bias update for the first ten and a signed bias for the last one. The results are presented in Fig. 4, which depicts the iterative procedure of BF-DCQO and how the obtained outcomes move towards the desired solution. The amount of resources needed per iteration is given in Table 2.
As a final remark, the final iteration of the purely quantum BF-DCQO routine exhibits a slightly worse performance when compared to the results of Fig. 3, which employed a noiseless simulation with a considerably lower gate threshold. The purely quantum BF-DCQO still outperforms the optimal solution obtained on the D-Wave platform, where we obtained $34.1\%$ of AR enhancement and $66.1\%$ of DS enhancement with respect the experiment run on D-Wave in Fig. 3.
@@
Table 2: 156-qubit NN HUBO results and resources needed on ibm_fez. We consider $n_{\text{shots}}=10000$, $\alpha=0.01$ and $\theta_{\text{cutoff}}=0.13$. At the last iteration, we performed signed bias.
@@
11 3 330 180 13 39 0.815 0.038
Table 3: 156-qubit NN MAX W-3-SAT HUBO results and resources needed on ibm_fez. We consider $n_{\text{shots}}=10000$, $\alpha=0.01$, and $\theta_{\text{cutoff}}=0.06$. In the last iteration, we apply a signed bias, where no entangling gates are used.
@@
Figure 6: 25-qubit dense MAX 3-SAT HUBO results on IonQ Aria 1 emulator. a-b, Cost function distributions for the first two and three iterations without noise (grey to maroon) and including an IonQ Aria 1 noise model profile plus debiased error mitigation (grey to teal), respectively. The exact ground state (continuous line) is reached for both scenarios. The dashed lines correspond to the maximum cost obtained from the distributions. The cost function is in units of the coupling $\omega_{ijk}$ in Eq. (16), which here is chosen as uniform. c-f, approximation ratios and distance to solution for both ideal and noisy emulations, respectively. Shaded areas indicate that the bias updating is changed from unsigned to signed bias at the last iteration. See Table 4.
Table 4: 25-qubit dense MAX 3-SAT HUBO results and resources needed on IonQ Aria 1 emulator. Noiseless (left) and IonQ Aria 1-noise profile with debiased error mitigation (right) simulations are considered with $n_{\text{shots}}=1000$, $\alpha=0.01$ and $\theta_{\text{cutoff}}=0.06$. At the last iteration, we performed signed bias, where no entangling gates were used.
Iteration $\text{MS}(\phi_{0},\phi_{1},\theta)$ $\text{GPi}(\phi)$ $\text{GPi2}(\phi)$ Depth AR DS
1 96 96 282 282 948 948 219 219 0.582 0.580 0.134 0.134
2 96 96 282 282 948 948 219 219 0.704 0.707 0.000 0.031
3 96 96 282 282 948 948 219 219 0.763 0.761 0.000 0.000
4 96 96 282 282 948 948 219 219 0.764 0.766 0.000 0.000
5 96 96 282 282 948 948 219 219 0.772 0.770 0.000 0.000
6 96 96 282 282 948 948 219 219 0.769 0.776 0.000 0.000
7 96 96 282 282 948 948 219 219 0.769 0.775 0.000 0.000
8 96 96 282 282 948 948 219 219 0.775 0.774 0.000 0.000
9 96 96 282 282 948 948 219 219 0.778 0.769 0.000 0.000
10 96 96 282 282 948 948 219 219 0.773 0.777 0.000 0.000
11 - - 26 26 52 52 7 7 0.979 0.943 0.000 0.000
@@
With the aim of solving more industrially relevant and denser large-scale problems, we also solve a 156-qubit NN MAX W-3-SAT instance on ibm_fez, described by Eq. (16) with only nearest-neighbour terms. We use the same parameters as for the previous instance but change $\theta_{\text{cutoff}}=0.06$ for convenience. The results are shown in Fig. 5, and the amount of resources needed per iteration is listed in Table 3. For this particular problem, the best outcome obtained is quite close to the optimal solution despite a purely quantum routine, which highlights the efficacy of BF-DCQO under noisy environments. Additionally, we solved the same HUBO instance on D-Wave using $n_{\text{shots}}=100000$. The purely quantum BF-DCQO again outperforms the D-Wave results: we obtain a $43.0\%$ enhancement of the AR and an $88.5\%$ enhancement of the DS with respect the experiment run on the D-Wave platform.
Dense problems with all-to-all connectivity
The sparsity and high locality of the terms involved in the instances considered on IBM render the problems amenable to classical simulation. To evaluate the efficacy of BF-DCQO in more intricate scenarios with long-range couplings, we extend our analysis to denser versions of MAX 3-SAT. In particular, we set as clause-to-variable ratio the critical density for 3-SAT, $\alpha_{C}=4.3$. Dense MAX 3-SAT instances with uniform couplings equal to unity are then generated randomly with the number of clauses given by $\lceil N\alpha_{C}\rceil$. It is important to note that the expected optimal solutions of MAX $k$-SAT instances, where clauses are uniformly weighted, can be either a product state or a low-entangled superposition of few bitstrings. As a last remark, optimal solution degeneracy was not present in our previous studies and it may compromise the performance.
We employ the 25-qubit IonQ Aria 1 emulator with native gate set given by $\{\text{MS}(\phi_{0},\phi_{1},\theta),\text{GPi}(\phi),\text{GPi2}(\phi)\}$ (see Supplementary Note 3). It mimics an all-to-all connected trapped-ion platform and encompasses both noiseless and noisy simulations, applying debiased error mitigation [101] with a noise profile that matches the Aria 1 expected performance. The system size of the problem allows to obtain the exact optimal solution to be obtained by our protocol. The results for $n_{\text{shots}}=1000$, $\alpha=0.01$ and $\theta_{\text{cutoff}}=0.06$ can be seen in Fig. 6, where the amount of resources required at each iteration presented in Table 4. The data shows that the BF-DCQO protocol is able to reach the exact optimal state in both noiseless and noisy environments at the second iteration, which serves to further validate the efficacy of the method.
Entanglement in the iterative optimization
The role of entanglement dynamics in quantum algorithms is an interesting and nontrivial question. In adiabatic quantum optimization, both the initial state and the desired final state are product states or low-entangled multipartite states, while the adiabatic path may explore highly entangled states [81, 80]. To gain insight into our observation that the required number of entangling gates is reduced with increasing iterations, we consider the flow of entanglement along the dynamical process.
We compute the entanglement entropy of Eq. (18) for the 433-qubit NN and the 25-qubit dense MAX 3-SAT HUBO instances solved previously in Figs. 2 and 6. As mentioned before, the first case corresponds to a sparse problem with a nondegenerate optimal solution and the second one to a dense problem with a degenerate optimal solution. For the second case, we only consider the noisy simulation including debiased error mitigation [Fig. 6b]. We show, in Fig. 7, how the entanglement entropy decays with the subsequent BF-DCQO iterations for both problems. Due to the dense nonlocal couplings, the average entanglement entropy generated in the second instance is larger than in the first. The numbers of entangling gates used here are listed in Tables 1 and 4. We can see that for the second instance, in the first three iterations, the average entanglement entropy decreases even when the number of entangling gates remains the same. Similar to the AR and DS in Figs. 2 and 6, the entropy then nearly saturates up to the last iteration. The signed bias applied in the last iteration significantly reduces the number of entangling gates and the entanglement entropy drops to a value close to zero. This indicates that the iteration protocol produces a classical ground state as expected.
Figure 7: Average entanglement entropy $S$ as a function of the iteration. The first instance corresponds to 433-qubit NN HUBO problem and the second instance to a 25-qubit dense MAX 3-SAT HUBO problem. Signed bias is applied at the last iteration, significantly reducing the number of entangling gates used and consequently the entanglement entropy. For the second instance, the entanglement entropy drops to zero at the last iteration as there are no entangling gates. We therefore leave out the last data point.
We note that while the entanglement entropy analyzed here is naturally obtained from the MPS representation, it is not a straightforward quantity to measure experimentally. In order to characterize entanglement in experiments run on quantum hardware, for instance in the case of highly entangled states not amenable to tensor network simulations, alternative measures would be required, such as correlator functions by means of quantum measurements [102, 90].
Conclusions